OpenStudy (anonymous):

solve the system of equations. 2x-y=8 y=x-3

HanAkoSolo (jamierox4ev3r):

2x-(x-3)=8 2x-x+3=8 x+3=8 -3 -3 x=5

OpenStudy (anonymous):

so y=2 (5,2)

HanAkoSolo (jamierox4ev3r):

y=5-3 y=2 2(5)-y=8 10-y=8 -10 -10 -y=-2 y=2

HanAkoSolo (jamierox4ev3r):

yes that is correct good job!

OpenStudy (anonymous):

thank you.